### v1

# Correcting Annotator Bias in Training Data: Population-Aligned Instance Replication (PAIR)

Abstract
Models trained on crowdsourced labels may not reflect broader population views when annotator pools are not representative. Since collecting representative labels is challenging, we propose Population-Aligned Instance Replication (PAIR), a method to address this bias through statistical adjustment. Using a simulation study of hate speech and offensive language detection, we create two types of annotators with different labeling tendencies and generate datasets with varying proportions of the types. Models trained on unbalanced annotator pools show poor calibration compared to those trained on representative data. However, PAIR, which duplicates labels from underrepresented annotator groups to match population proportions, significantly reduces bias without requiring new data collection. These results suggest statistical techniques from survey research can help align model training with target populations even when representative annotator pools are unavailable. We conclude with three practical recommendations for improving training data quality.  [^1] The code for experiments will be available at https://github.com/soda-lmu/PAIR.
Correcting Annotator Bias in Training Data:
Population-Aligned Instance Replication (PAIR)
[…]
1 Introduction and Inspiration
NLP models should align with the interests and judgments of the population they impact Sorensen et al. (2024); Fleisig et al. (2024). However, the training and feedback data for these models often comes from crowdworkers or convenience samples of annotators (e.g.  student assistants). These populations differ from the general population on important characteristics like age, education, and cultural context Smart et al. (2024); Berinsky et al. (2012); Ouyang et al. (2022), and these characteristics impact the labels they assign Sap et al. (2022); Fleisig et al. (2023); Kirk et al. (2024).
Figure 1: Top: Adjusting survey data to match population produces high quality results. Bottom: Can a similar adjustment in data annotations also improve model performance?
Fortunately, survey researchers have developed robust statistical techniques to estimate population-level parameters from non-representative samples Eckman et al. (2024); Bethlehem et al. (2011a). The top panel of Figure 1 shows a simple workflow of collecting survey data and then creating statistical weights to match the data to the population. We propose that similar techniques could help align machine learning models with target populations, even when working with imperfect annotator pools (bottom panel).
To test this approach, we use a simulation study, a common approach in the statistics literature (Burton et al., 2006; Valliant, 2019; Morris et al., 2019). We simulate seven populations, create several labeled datasets with varying mixes of annotators, and train model on each dataset. We investigate two research questions: RQ1: How does the composition of the annotator pool impact model calibration and performance? RQ2: Can techniques from survey methodology mitigate the effects of non-representative annotator pools?
Our results demonstrate that models trained on nonrepresentative annotator pools perform poorly. However, simple adjustment methods can improve performance without collecting additional data. These findings suggest that insights from survey methodology help make AI systems more representative of the populations they serve.
2 Annotation Simulation
To address our research questions, we create populations with two types of people: those more likely to perceive offensive language and hate speech and those less likely. We then simulate labels, varying the mix of the two types of people, to approximate when happens when the pool of annotators does not represent the characteristics of the population.
Data.
We use a dataset [^2] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2 of 3,000 tweets sampled from Davidson et al. (2017). Each tweet has 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no) (Kern et al., 2023). We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with.
We randomly select (without replacement) 12 labels (of both OL and HS) from the 15 labels of each tweet in the original dataset. [^3] For reasons that will become clear, it is helpful for the number of labels per tweet to be divisible by four. Let $p_{i,OL}$ be the proportion of the 12 annotators who labeled tweet $i$ as OL and $p_{i,HS}$ defined similarly. Table 1 shows the distribution of these proportions.
Variable p25 Median Mean p75
$p_{i,OL}$ 0.167 0.667 0.564 0.917
$p_{i,HS}$ 0.083 0.167 0.301 0.50
p refers to percentile
Table 1: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in Gold Dataset
Population Setup.
We imagine populations made up of equal shares of two types of people. Type A people are less likely to say a tweet contains OL or HS. Type B people are more likely:
$p_{i,OL}^{A}$ $=\text{max}(p_{i,OL}-\beta,0)$ (1)
$p_{i,OL}^{B}$ $=\text{min}(p_{i,OL}+\beta,1)$ (2)
The HS probabilities are defined similarly. Here $\beta$ captures the magnitude of the bias. We vary $\beta$ from $[0.05,0.3]$ by $0.05$ to create seven populations and seven vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B},p_{i,HS}^{A},p_{i,HS}^{B}$) for each tweet.
Label Simulation.
For each value of $\beta$, we create four datasets, each with 3,000 tweets (Table 2). The Balanced Dataset contains OL labels from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$). The proportion of A and B annotators in this dataset matches the population. The labels in this dataset are our gold standard.
Dataset Labels per tweet A labels B labels
Balanced 12 6 6
Unbalanced 1 9 6 3
Unbalanced 2 12 9 3
Adjusted 12 6 3 + 3*
NA = Not Applicable; * 3 B labels duplicated
Table 2: Four Training Datasets for each label (OL, HS) and each bias value ($\beta$)
We then create two unbalanced datasets. Unbalanced 1 randomly deletes three B labels for each tweet from the Balanced Dataset. Unbalanced 2 adds three additional A labels, drawn from $p_{i}^{A}$, to the Unbalanced 1 dataset. The Unbalanced 2 Dataset is more unbalanced than Unbalanced 1, but contains the same number of annotations as the Balanced dataset.
Finally, we create the Adjusted Dataset. It is the same as the Unbalanced 1 dataset, but the B annotations are duplicated. This duplication is an easy way to adjust the unbalanced training dataset to reflect the population; a generalized version of this adjustment is provided in A.5.
Appendix Figure 5 shows the percentage of tweets labeled OL and HS in the four datasets for each value of $\beta$.
3 Model Training and Evaluation
Training and Test Setup.
We train models on each dataset. We divide each dataset, at the tweet level, into training (2000 tweets), development (500), and test (500) sets.
Model Selection and Training.
We used RoBERTa base Liu et al. (2019) as our text classifier. The model trained for five epochs on each dataset, with development set optimization. To ensure reliable results, we trained five versions with different random seeds and averaged their performance. Appendix §A.2 contains additional details.
Performance Metrics.
We measure model performance using two metrics.
  • Absolute Calibration Bias (ACB): This metric compares the observed proportion of OL/HS labels in the Balanced dataset to each the model’s predicted probabilities ($\text{preds}_{i}$):
    $\text{ACB}_{OL}=\frac{1}{n}\sum_{i=1}^{n}\left|p_{i,OL}-\text{preds}_{i,OL}\right|$ (3)
    ACB simplifies the ECE metric Naeini et al. (2015); Guo et al. (2017) by removing bins, and makes the single difference more straightforward by incorporating true frequencies directly. A low ACB score means the model’s predicted probabilities match the true frequencies in the Gold dataset.
  • F1 Score: This standard classification metric balances precision and recall.
The appendix also presents ECE (similar to ACB) and ROC-AUC (similar to F1) metrics (see §A.4). We report all metrics on the test set.
4 Results
Model Calibrations.
Figure 2 compares the ACB in the test set for models trained on the simulated datasets. Lower ACB indicates better model calibrations with the test labels. The dark lines show average ACB across the five training runs and the shading shows the standard deviation. See Appendix Figure 7 for the ECE results, which are similar.
In the OL graph, ACB for the models trained on the Balanced and Adjusted datasets do not increase with $\beta$ and are close together. ACB for the models trained on the two unbalanced datasets is greater and grows with $\beta$. These results demonstrate the effectiveness of our adjustment method. Duplicating the labels from the underrepresented annotator type to match population proportions improves calibration.
In the HS graph, the trends are less clear. Because HS is rarer in the dataset (Table 1), the A annotations are often 0, which complicates interpretation of the HS panel of Figure 2. We address this issue with additional analyses below.
Figure 2: Model ACB scores, by dataset and bias ($\beta$)
Model Predictions.
Figure 3 compares the models’ F1 scores. In contrast to Figure 2, we do not see strong differences between the models trained on the different datasets. For all datasets, model performance declines with $\beta$: as the amount of noise in the labels increases, the models have a harder time predicting the labels. The ROC-AUC graphs show the same pattern (Appendix Figure 8).
Because the F1 metric focuses on binary predictions, it is less sensitive to training biases compared to calibration metrics, which more explicitly capture biases through prediction scores. These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models, and binary classification performance alone can obscure such effects. In decision-making, miscalibrated predictions can have harmful consequences when, e.g., hateful content remains undetected Van Calster et al. (2019).
Figure 3: Model F1 scores, by dataset and bias ($\beta$)
Results with Difficult Tweets.
Our simulations assumed that all tweets are impacted the same way (Eq. (2)), which is an oversimplification. More realistically, annotator characteristics likely have more impact for ambiguous tweets. We repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$. This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq. (2)). The filtered datasets contain 267 (OL) and 360 (HS) tweets.
Figure 4 shows the ACB for models trained on the filtered datasets. In the OL graph, the patterns in the ACB scores are similar to Figure 2. The Balanced and Adjusted models have similar ACB and are lower than the Unbalanced models. The HS graph shows the same pattern, which is in contrast to Figure 2 where HS performed differently.
The F1 results for this subset of tweets are in Appendix Figure 6. The OL graph shows higher F1 scores for the Balanced and Adjusted models than for the two unbalanced models. There is no clear pattern in the F1 scores for the F1 model, though we do see an unexpected decrease in agreement in the Adjusted model when $\beta=0.3$.
Figure 4: Model ACB scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$), by dataset and bias ($\beta$)
5 Discussion & Recommendations
Our results demonstrate two key findings about annotator representation in training data. First, models perform less well when trained on data from non-representative annotator pools. Second, simple statistical adjustments can help correct for these biases without collecting additional data. These findings have important implications for dataset creation and model training.
Recommendations
If these findings hold up in future work, we recommend the AI/ML field take three steps:
1) Identify which annotator characteristics influence labeling and feedback decisions for different tasks. Research should integrate social science methods to understand how demographics, attitudes, and behaviors shape annotation patterns Eckman et al. (2024). This knowledge would help determine which characteristics need to be balanced across annotator pools for specific tasks.
2) Begin collecting relevant characteristics from annotators and gather corresponding population-level data. National censuses and large-scale surveys could provide useful population benchmarks. [^4] Collection and release of annotator characteristics or weights derived from them may raise concerns about annotator confidentiality. This topic is outside the scope of this paper, however, the survey literature contains useful approaches (see Karr, 2016, for a review). We also note that collecting annotator characteristics may require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al. (2024).
3) Implement data adjustment methods that account for differences between annotator and population characteristics. While our simple duplication approach showed promise, more sophisticated statistical techniques from survey research may yield better results.
Limitations
Stylized Biases.
Our simulation makes strong assumptions about annotator behavior, particularly in modeling consistent biases across annotator types. Real-world annotator biases may be more nuanced or context-dependent. Future work could incorporate more realistic biases and refine the proposed simulations and statistical techniques.
Sampling Variability.
We have created only one version the four datasets for each label type and value of $\beta$, each of which contains random draws from the Bernoulli distribution. A more traditional statistical approach would create multiple versions of the datasets and train models on each one, to average over the sampling variability. We have not done that in this preliminary study because of the high cost and time needed to fine tune many RoBERTa models. As discussed, we have used five seeds in model training.
Generalization Beyond Task Types.
The study focuses only on binary classification tasks. Many real-world annotation tasks involve multiple classes or labels, which may show different bias patterns. Additional research is needed to extend these methods to more complex classification scenarios.
Evaluation Metrics.
While we measured calibration and classification accuracy, we did not examine other important metrics such as fairness across subgroups or robustness to adversarial examples. Future work should on training data adjustment should assess a broader range of performance measures.
Ethical Considerations
In this simulation study, we experiment on a publicly available dataset Kern et al. (2023), which contains offensive and hateful tweets. We do not support the views expressed in the tweets. The simulation study itself does not collect any new data or raise any ethical considerations.
Acknowledgments
We acknowledge use of the Claude model to edit the text of the paper and to assist in coding. We thank the members of SODA Lab and MaiNLP labs from LMU Munich, and the members of the Social Data Science Group from University of Mannheim for their constructive feedback. This research is partially supported by RTI International, MCML, and BERD@NFDI. BP is supported by ERC Consolidator Grant DIALECT (101043235).
Appendix A Appendix
A.1 Data
Figure 5 shows the percentage of tweets labeled OL and HS in the four datasets for each value of $\beta$. The percentage in the Adjusted dataset is similar to that in the Balanced dataset for all values of $\beta$. The two unbalanced datasets have lower rates of OL and HS, because they overrepresent the A annotators, who are less likely to label OL and HS.
Figure 5: Percentage of OL, HS “yes” labels, by dataset and bias ($\beta$)
Because HS is rare in our dataset, as $\beta$ increases, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ values increase. This issue leads to a higher proportion of “yes” HS labels in the Balanced and Adjusted datasets, which have more B labels than the unadjusted datasets. [^5] If we had picked the B labels to overrepresent, we would have seen a converse problem with the OL labels.
A.2 Model Training Details
Our implementation of RoBERTa models was based on the libraries pytorch (Paszke et al., 2019) and transformers (Wolf et al., 2020). During training, we used the same hyperparameter settings of the respective models for our five training conditions to keep these variables consistent for comparison purposes. We report the hyperparameter settings of the models in Table 3. To avoid random effects on training, we trained each model variation with five random seeds $\{10,42,512,1010,3344\}$ and took the average across the models. All experiments were conducted on an NVIDIA® A100 80 GB RAM GPU.
Hyperparameter Value
encoder roberta-base
epochs_trained $5$
learning_rate $3e^{-5}$
batch_size $32$
warmup_steps $500$
optimizer AdamW
max_length $128$
Table 3: Hyperparameter settings of RoBERTa models
A.3 Additional Results
We present the results of F1 scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$) in Figure 6.
Figure 6: Model F1 scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$), by dataset and bias ($\beta$)
A.4 Additional Metrics
We also use the Expected Calibration Error (ECE) metric to evaluate model calibration. This metric quantifies calibration quality by measuring the weighted average absolute difference between accuracy and predicted confidence across $M$ bins:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\cdot\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (4)
where $B_{m}$ is the set of samples in the $m$-th bin, $|B_{m}|$ its size, $n$ the total number of samples, $\text{acc}(B_{m})$ the accuracy, and $\text{conf}(B_{m})$ the mean predicted confidence in $B_{m}$.
Figure 7 shows the results. Consistent with our main findings in §4, we observe that for OL labels, the ECE values are relatively stable on the Balanced and Adjusted datasets ($t=-0.48,p=0.64$), with divergence with $\beta=0.15$. This contrasts with the significant divergence in ECE values for the Unbalanced subsets as the $\beta$ value increases.
Figure 7: Model ECE metrics, by dataset and bias ($\beta$)
Additionally, for model performance, we use another metric ROC-AUC in Figure 8. The results are very similar to the F1 results in §4.
Figure 8: Model ROC-AUC metrics, by dataset and bias ($\beta$)
A.5 Adjustment Details
The adjustment we use to make the labeler pool more representative of the target population is a form of pseudo-population generation (Quatember, 2015). We create the pseudo-population by first constructing post-stratification weights, performing weight normalization to ensure the sum of the weights equals the size of the target population, and then duplicating each observation proportionally to its weight via deterministic replication.
Post-stratification Bethlehem et al. (2011b) is a method of statistical adjustment that makes a selected sample more closely resemble a target population. Post-stratification requires population level totals or proportions and corresponding case-level observations of the same characteristic. to be available for each sample stratum that will be used in the weighting. The weight for unit $i$ is determined for each stratum by:
$w_{i}=\frac{P_{s}}{S_{s}}$ (5)
where $P_{s}$ is the true population proportion (or total) for stratum (or group) $s$ and $S_{s}$ is the sample proportion (or total) for stratum $s$. In our case, the strata of interest was a single variable (labeler Types A & B). However, post-stratification can involve multiple variables if their joint distribution is known at the population level.
Although the post-stratified weights will preserve the ratios of the strata in the target population, the weighted totals themselves may not match those in the target population. Weight normalization can be used to address this by updating the survey weights so that they sum to a desired total. The normalized weight for unit $i$ can be calculated by:
$w_{i}^{\text{normalized}}=w_{i}^{\text{initial}}\cdot\frac{T}{\sum_{i=1}^{n}w_{i}^{\text{initial}}}$ (6)
where $T$ is the target total. Since we want the Adjusted dataset to match the size of our gold standard Balanced dataset, the target total for the simulation was 12 labels per tweet.
Lastly, to construct a pseudo-population from our weighted data, we perform deterministic replication by replicating each unit $n_{i}$ times. In this initial work, we appreciated the simple interpretation and reproducibility of this approach. However, researcher may prefer other approaches, such as replication via resampling, if they are interested in how the adjustment varies across samples.
In our case, after post-stratification and weight normalization, each Type A label in the Unbalanced 1 dataset receives a weight of 1 and each Type B label receives a weight of 2. This resulted in an Adjusted dataset where each of the Type A labels stays the same and each of the Type B labels is duplicated once.

### v2

@@
 Abstract
-Models trained on crowdsourced labels may not reflect broader population views when annotator pools are not representative.
-Since collecting representative labels is challenging, we propose Population-Aligned Instance Replication (PAIR), a method to address this bias through statistical adjustment.
-Using a simulation study of hate speech and offensive language detection, we create two types of annotators with different labeling tendencies and generate datasets with varying proportions of the types.
-Models trained on unbalanced annotator pools show poor calibration compared to those trained on representative data.
-However, PAIR, which duplicates labels from underrepresented annotator groups to match population proportions, significantly reduces bias without requiring new data collection.
-These results suggest statistical techniques from survey research can help align model training with target populations even when representative annotator pools are unavailable.
-We conclude with three practical recommendations for improving training data quality.
-[^1] The code for experiments will be available at https://github.com/soda-lmu/PAIR.
+Models trained on crowdsourced labels may not reflect broader population views, because those who work as annotators do not represent the population.
+We propose Population-Aligned Instance Replication (PAIR), a method to address bias caused by non-representative annotator pools.
+Using a simulation study of offensive language and hate speech, we create two types of annotators with different labeling tendencies and generate datasets with varying proportions of the types.
+We observe that models trained on unbalanced annotator pools show poor calibration compared to those trained on representative data.
+By duplicating labels from underrepresented annotator groups to match population proportions, PAIR reduces bias without collecting additional annotations.
+These results suggest that statistical techniques from survey research can improve model performance.
+We conclude with practical recommendations for improving the representativity of training data and model performance.
+[^1] The code for experiments is available at https://github.com/soda-lmu/PAIR.
 Correcting Annotator Bias in Training Data:
@@
 1 Introduction and Inspiration
-NLP models should align with the interests and judgments of the population they impact Sorensen et al.
+When a hate speech detection model flags harmless expressions as toxic, or a content moderation system fails to identify genuinely harmful content, the root cause often lies not in the model architecture, but in who labeled the training data.
+While NLP models aim to serve broad populations, the human judgments used to train these systems typically come from a non-representative pool of annotators – crowdworkers and convenience samples whose demographics, cultural contexts, and worldviews may differ from the communities the models ultimately impact Sorensen et al.
 (2024); Fleisig et al.
 (2024).
-However, the training and feedback data for these models often comes from crowdworkers or convenience samples of annotators (e.g.
-student assistants).
-These populations differ from the general population on important characteristics like age, education, and cultural context Smart et al.
-(2024); Berinsky et al.
-(2012); Ouyang et al.
-(2022), and these characteristics impact the labels they assign Sap et al.
+These non-representative annotator pools can have real consequences, because annotator characteristics like age, education level, and cultural background impact how content is labeled Sap et al.
 (2022); Fleisig et al.
@@
 (2024).
+Models trained on non-representative data can perpetuate the biases and blind spots of their training data Smart et al.
+(2024); Berinsky et al.
+(2012); Ouyang et al.
+(2022); Mehrabi et al.
+(2021); Rolf et al.
+(2021); Favier et al.
+(2023); Hebert-Johnson et al.
+(2018); Hüllermeier and Waegeman (2021).
 Figure 1: Top: Adjusting survey data to match population produces high quality results. Bottom: Can a similar adjustment in data annotations also improve model performance?
-Fortunately, survey researchers have developed robust statistical techniques to estimate population-level parameters from non-representative samples Eckman et al.
-(2024); Bethlehem et al.
-(2011a).
-The top panel of Figure 1 shows a simple workflow of collecting survey data and then creating statistical weights to match the data to the population.
-We propose that similar techniques could help align machine learning models with target populations, even when working with imperfect annotator pools (bottom panel).
-To test this approach, we use a simulation study, a common approach in the statistics literature (Burton et al., 2006; Valliant, 2019; Morris et al., 2019).
-We simulate seven populations, create several labeled datasets with varying mixes of annotators, and train model on each dataset.
-We investigate two research questions: RQ1: How does the composition of the annotator pool impact model calibration and performance?
-RQ2: Can techniques from survey methodology mitigate the effects of non-representative annotator pools?
+Fortunately, survey researchers have developed robust statistical techniques to produce population-level estimates from non-representative samples Bethlehem et al.
+(2011).
+The top panel of Figure 1 shows a simple survey workflow: collecting survey data, creating statistical weights to match the sample to the population, and estimating population parameters.
+We propose adapting this approach to the machine learning context, enabling models to better align with target populations even when trained on non-representative annotator pools (bottom panel).
+To test this approach, we conduct a rigorous simulation study following established practices in statistical research (Burton et al., 2006; Valliant, 2019; Morris et al., 2019).
+Using controlled experiments, simulated populations, and multiple training datasets varying in annotator composition, we investigate two questions:
+  • RQ1: How do non-representative annotator pools impact model calibration and accuracy?
+  • RQ2: Can survey techniques effectively mitigate these annotator pool effects?
 Our results demonstrate that models trained on nonrepresentative annotator pools perform poorly.
 However, simple adjustment methods can improve performance without collecting additional data.
-These findings suggest that insights from survey methodology help make AI systems more representative of the populations they serve.
-2 Annotation Simulation
-To address our research questions, we create populations with two types of people: those more likely to perceive offensive language and hate speech and those less likely.
-We then simulate labels, varying the mix of the two types of people, to approximate when happens when the pool of annotators does not represent the characteristics of the population.
-Data.
-We use a dataset [^2] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2 of 3,000 tweets sampled from Davidson et al.
-(2017).
-Each tweet has 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no) (Kern et al., 2023).
-We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with.
-We randomly select (without replacement) 12 labels (of both OL and HS) from the 15 labels of each tweet in the original dataset.
-[^3] For reasons that will become clear, it is helpful for the number of labels per tweet to be divisible by four.
+These findings suggest that insights from survey methodology can help AI systems better represent the populations they serve.
+2 Related Work
+Several strands of related work inform our approach to identifying and mitigating bias due to the use of non-representative annotators:
+Annotator Impact on Data and Models.
+Annotator characteristics and attitudes significantly influence label quality, particularly for subjective tasks like toxicity detection (Giorgi et al., 2024; Prabhakaran et al., 2021; Fleisig et al., 2023; Sap et al., 2022).
+For example, annotators’ political views and racial attitudes affect their toxicity judgments (Sap et al., 2022).
+Models trained on non-representative annotator pools inherit these biases and generalize poorly (Smart et al., 2024; Berinsky et al., 2012; Ouyang et al., 2022; Mehrabi et al., 2021; Rolf et al., 2021; Favier et al., 2023).
+Annotator Demographics.
+Several researchers advocate collecting annotator demographics to assess representation and identify biases (Bender and Friedman, 2018; Prabhakaran et al., 2021).
+[^2] In our context, these characteristics are used only to analyze bias; because they are not available for unlabeled text, they are not features that the model can use.
+However, collecting and releasing these data can raise privacy concerns (Fleisig et al., 2023).
+Debiasing & Data Augmentation Methods.
+Prior work has proposed various approaches to reduce bias in training data features and labels (Calders et al., 2009; Kamiran and Calders, 2012; Sharma et al., 2020).
+Most similar to our work is the resampling and reweighting approaches of Calders et al.
+(2009) and Kamiran and Calders (2012), and the oversampling of minority class cases of Ling and Li (1998).
+PAIR adapts these methods to balance annotator characteristics rather than class labels or sensitive observation-level features.
+PAIR retains the simplicity and interpretability of earlier resampling methods while extending them to a “Learning with Disagreement” (Uma et al., 2021; Leonardelli et al., 2023) setting with multiple annotations per observation, by replicating labels from underrepresented annotator groups.
+3 Annotation Simulation and Model Training
+To address our research questions, we imagine a population made up of two types of people: those more likely to perceive offensive language and hate speech and those less likely.
+We create three sets of simulated annotations which differ in the mix of the annotator types.
+We then create a fourth data set, using the PAIR algorithm, to attempt to fix the imbalance in the annotators.
+We fine-tune RoBERTa models on the four data sets and evaluate the effect of annotator characteristics on model performance (RQ1) and the ability of the PAIR algorithm to improve performance (RQ2).
+3.1 Simulating Annotations
+We use a dataset of 3,000 English-language tweets, each with 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no) (Kern et al., 2023).
+[^3] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2 We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with.
+We randomly select (without replacement) 12 labels (of both OL and HS) of each tweet in the original dataset.
+[^4] As shown in Table 2, we can more carefully control the construction of our data sets when the number of labels per tweet is divisible by four.
 Let $p_{i,OL}$ be the proportion of the 12 annotators who labeled tweet $i$ as OL and $p_{i,HS}$ defined similarly.
-Table 1 shows the distribution of these proportions.
-Variable p25 Median Mean p75
+Table 1 shows the distribution of these proportions across the 3,000 tweets.
+The HS labels are unevenly distributed, whereas the OL labels are relatively balanced.
+Because this work is a preliminary investigation of the PAIR approach, and balanced labels provide a clearer view of PAIR’s abilities, the main body of the paper focuses on results with the OL labels; the HS results in Appendix §C.
+Variable $25^{th}$ pctile. Median Mean $75^{th}$ pctile.
 $p_{i,OL}$ 0.167 0.667 0.564 0.917
 $p_{i,HS}$ 0.083 0.167 0.301 0.50
-p refers to percentile
-Table 1: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in Gold Dataset
-Population Setup.
-We imagine populations made up of equal shares of two types of people. Type A people are less likely to say a tweet contains OL or HS. Type B people are more likely:
+Table 1: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in original data
+We imagine a population made up of equal shares of two types of people. Type A people are less likely to say a tweet contains OL. Type B people are more likely:
 $p_{i,OL}^{A}$ $=\text{max}(p_{i,OL}-\beta,0)$ (1)
 $p_{i,OL}^{B}$ $=\text{min}(p_{i,OL}+\beta,1)$ (2)
-The HS probabilities are defined similarly.
 Here $\beta$ captures the magnitude of the bias.
-We vary $\beta$ from $[0.05,0.3]$ by $0.05$ to create seven populations and seven vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B},p_{i,HS}^{A},p_{i,HS}^{B}$) for each tweet.
-Label Simulation.
-For each value of $\beta$, we create four datasets, each with 3,000 tweets (Table 2).
-The Balanced Dataset contains OL labels from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$).
-The proportion of A and B annotators in this dataset matches the population.
-The labels in this dataset are our gold standard.
+We vary $\beta$ from $[0.05,0.3]$ by $0.05$, corresponding to an increase or decrease in the probability to judge a tweet as OL by five to 30 percentage points.
+This range is large on the probability scale and covers most reasonable situations.
+With these seven values of $\beta$, we create seven vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B}$) for each tweet.
+We then create four datasets, each with 3,000 tweets (Table 2), for each value of $\beta$.
+The Representative Dataset contains OL labels from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$).
+Because the proportion of A and B annotators in this dataset matches the population, the labels in this dataset are our gold standard.
 Dataset Labels per tweet A labels B labels
-Balanced 12 6 6
-Unbalanced 1 9 6 3
-Unbalanced 2 12 9 3
+Representative 12 6 6
+Non-representative 1 9 6 3
+Non-representative 2 12 9 3
 Adjusted 12 6 3 + 3*
-NA = Not Applicable; * 3 B labels duplicated
-Table 2: Four Training Datasets for each label (OL, HS) and each bias value ($\beta$)
+* 3 B labels duplicated
+Table 2: Four training datasets for each bias value ($\beta$)
 We then create two unbalanced datasets.
-Unbalanced 1 randomly deletes three B labels for each tweet from the Balanced Dataset.
-Unbalanced 2 adds three additional A labels, drawn from $p_{i}^{A}$, to the Unbalanced 1 dataset.
-The Unbalanced 2 Dataset is more unbalanced than Unbalanced 1, but contains the same number of annotations as the Balanced dataset.
-Finally, we create the Adjusted Dataset.
-It is the same as the Unbalanced 1 dataset, but the B annotations are duplicated.
-This duplication is an easy way to adjust the unbalanced training dataset to reflect the population; a generalized version of this adjustment is provided in A.5.
-Appendix Figure 5 shows the percentage of tweets labeled OL and HS in the four datasets for each value of $\beta$.
-3 Model Training and Evaluation
+Non-representative 1 randomly deletes three B labels for each tweet from the Representative Dataset.
+Non-representative 2 adds three additional A labels, drawn from $p_{i}^{A}$, to the Non-representative 1 dataset.
+The Non-representative 2 Dataset is more unbalanced than Non-representative 1, but contains the same number of annotations as the Representative dataset.
+Finally, we use the PAIR algorithm to create the Adjusted Dataset.
+It is the same as the Non-representative 1 dataset, but the three B annotations are duplicated.
+This duplication is an easy way to adjust the unbalanced training dataset to reflect the population.
+Appendix §A provides a general version of the PAIR algorithm which can handle imbalances across multiple annotator characteristics.
+Figure 2 shows the percentage of tweets labeled OL in the four datasets for each value of $\beta$.
+Figure 2: Percentage of tweets annotated as OL, by dataset and bias ($\beta$)
+3.2 Model Training and Evaluation
 Training and Test Setup.
-We train models on each dataset. We divide each dataset, at the tweet level, into training (2000 tweets), development (500), and test (500) sets.
+We train models on each dataset.
+We divide each dataset, at the tweet level, into training (2000 tweets), development (500), and test (500) sets.
+Each tweet appears 12 times in the Representative, Non-representative 2, and Adjusted data sets and nine times in the Non-representative 1 set.
 Model Selection and Training.
 We used RoBERTa base Liu et al.
-(2019) as our text classifier.
-The model trained for five epochs on each dataset, with development set optimization.
+(2019) as our text classifier, training for five epochs on each dataset, with development set optimization.
 To ensure reliable results, we trained five versions with different random seeds and averaged their performance.
-Appendix §A.2 contains additional details.
+Appendix §B contains details on model training.
 Performance Metrics.
-We measure model performance using two metrics.
-  • Absolute Calibration Bias (ACB): This metric compares the observed proportion of OL/HS labels in the Balanced dataset to each the model’s predicted probabilities ($\text{preds}_{i}$):
+We evaluate models using both calibration and accuracy metrics on the test set.
+While accuracy metrics directly measure classification performance, calibration metrics provide crucial insights into model reliability by assessing probability estimate quality – particularly important for high-stakes applications requiring trustworthy confidence measures.
+For calibration, we report Absolute Calibration Bias (ACB), which measures how well model predicted probabilities match true frequencies in the Representative dataset (lower is better).
 $\text{ACB}_{OL}=\frac{1}{n}\sum_{i=1}^{n}\left|p_{i,OL}-\text{preds}_{i,OL}\right|$ (3)
-    ACB simplifies the ECE metric Naeini et al.
-(2015); Guo et al.
-(2017) by removing bins, and makes the single difference more straightforward by incorporating true frequencies directly.
-A low ACB score means the model’s predicted probabilities match the true frequencies in the Gold dataset.
-  • F1 Score: This standard classification metric balances precision and recall.
-The appendix also presents ECE (similar to ACB) and ROC-AUC (similar to F1) metrics (see §A.4). We report all metrics on the test set.
+For accuracy, we report the F1 score.
 4 Results
-Model Calibrations.
-Figure 2 compares the ACB in the test set for models trained on the simulated datasets.
-Lower ACB indicates better model calibrations with the test labels.
-The dark lines show average ACB across the five training runs and the shading shows the standard deviation.
-See Appendix Figure 7 for the ECE results, which are similar.
-In the OL graph, ACB for the models trained on the Balanced and Adjusted datasets do not increase with $\beta$ and are close together.
+Calibration.
+Figure 3(1) compares the ACB in the test set for models trained on the simulated datasets. The dark lines show average ACB across the five training runs and the shading shows the standard deviation.
+The ACB for the models trained on the Adjusted dataset closely tracks that for the Representative (gold standard) data set and does not increase with $\beta$.
 ACB for the models trained on the two unbalanced datasets is greater and grows with $\beta$.
 These results demonstrate the effectiveness of our adjustment method.
-Duplicating the labels from the underrepresented annotator type to match population proportions improves calibration.
-In the HS graph, the trends are less clear.
-Because HS is rarer in the dataset (Table 1), the A annotations are often 0, which complicates interpretation of the HS panel of Figure 2.
-We address this issue with additional analyses below.
-Figure 2: Model ACB scores, by dataset and bias ($\beta$)
-Model Predictions.
-Figure 3 compares the models’ F1 scores.
-In contrast to Figure 2, we do not see strong differences between the models trained on the different datasets.
-For all datasets, model performance declines with $\beta$: as the amount of noise in the labels increases, the models have a harder time predicting the labels.
-The ROC-AUC graphs show the same pattern (Appendix Figure 8).
-Because the F1 metric focuses on binary predictions, it is less sensitive to training biases compared to calibration metrics, which more explicitly capture biases through prediction scores.
-These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models, and binary classification performance alone can obscure such effects.
-In decision-making, miscalibrated predictions can have harmful consequences when, e.g., hateful content remains undetected Van Calster et al.
+Duplicating the labels from the underrepresented annotator type to match population proportions improves model calibration.
+Figure 3: Metrics for OL models, by dataset and bias ($\beta$)
+Accuracy.
+Figure 3(2) compares the models’ F1 scores.
+In contrast to Figure 3(1), we do not see strong differences between the models trained on the different datasets.
+For all datasets, model performance declines with $\beta$: as the amount of bias in the labels increases, the models have a harder time predicting the binary OL label.
+Because the F1 metric focuses on binary predictions, it is less sensitive to training biases than calibration metrics like ACB, which more explicitly capture biases through prediction scores.
+These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models: binary classification metrics can obscure such effects.
+In decision-making, miscalibrated predictions can have harmful consequences when, for example, hateful content remains undetected Van Calster et al.
 (2019).
-Figure 3: Model F1 scores, by dataset and bias ($\beta$)
-Results with Difficult Tweets.
-Our simulations assumed that all tweets are impacted the same way (Eq.
-(2)), which is an oversimplification.
-More realistically, annotator characteristics likely have more impact for ambiguous tweets.
-We repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$.
-This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq.
-(2)).
-The filtered datasets contain 267 (OL) and 360 (HS) tweets.
-Figure 4 shows the ACB for models trained on the filtered datasets.
-In the OL graph, the patterns in the ACB scores are similar to Figure 2.
-The Balanced and Adjusted models have similar ACB and are lower than the Unbalanced models.
-The HS graph shows the same pattern, which is in contrast to Figure 2 where HS performed differently.
-The F1 results for this subset of tweets are in Appendix Figure 6.
-The OL graph shows higher F1 scores for the Balanced and Adjusted models than for the two unbalanced models.
-There is no clear pattern in the F1 scores for the F1 model, though we do see an unexpected decrease in agreement in the Adjusted model when $\beta=0.3$.
-Figure 4: Model ACB scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$), by dataset and bias ($\beta$)
+Ablation Study 1: Hate Speech labels.
+Appendix §C gives the ACB and F1 results for the HS labels.
+The PAIR approach has limited effectiveness when label distributions are highly skewed.
+This failure is likely due to the rarity of hate speech labels in the dataset (only 16.7% positive labels) combined with our simulation setup, which further reduces positive labels in the unbalanced datasets.
+PAIR may need modifications, such as stratified sampling or adaptive replication, when working with highly imbalanced datasets.
+Ablation Study 2: Difficult Tweets.
+A representative pool of annotators may be more important for instances that are difficult or about which people disagree.
+Filtering our tweets in this way also removes cases where floor and ceiling effects could mask the impact of annotator characteristics.
+To test PAIR performance on difficult tweets, we reran the models on tweets where $0.60.4<p_{i,HS}<$ 0.6, $0.60.4<p_{i,OL}<$ 0.6.
+Performance on OL is still good, and performance on HS is improved (Appendix §D).
 5 Discussion & Recommendations
-Our results demonstrate two key findings about annotator representation in training data.
-First, models perform less well when trained on data from non-representative annotator pools.
-Second, simple statistical adjustments can help correct for these biases without collecting additional data.
-These findings have important implications for dataset creation and model training.
-Recommendations
-If these findings hold up in future work, we recommend the AI/ML field take three steps:
-1) Identify which annotator characteristics influence labeling and feedback decisions for different tasks.
-Research should integrate social science methods to understand how demographics, attitudes, and behaviors shape annotation patterns Eckman et al.
+Our results show that (RQ1) OL prediction models perform less well when trained on data from non-representative annotator pools, and (RQ2) simple statistical adjustments can improve model calibration without collecting additional annotations.
+These findings establish a promising bridge between survey statistics and machine learning - offering a practical approach to make AI systems more representative of and responsive to the populations they serve, particularly for tasks involving subjective human judgments.
+We recommend these four steps to reduce bias due to non-representative annotator pools:
+  1. 1)
+    Use social science research to identify the annotator characteristics that influence the propensity to engage in annotation and the annotations provided Eckman et al. (2024).
+  2. 2)
+    Collect these characteristics from annotators and gathering corresponding population-level data from national censuses or high-quality surveys.
+[^5] Collection and release of annotator characteristics or weights derived from them may raise confidentiality concerns.
+The survey literature contains relevant approaches (see Karr, 2016, for a review).
+Collecting annotator characteristics may also require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al.
 (2024).
-This knowledge would help determine which characteristics need to be balanced across annotator pools for specific tasks.
-2) Begin collecting relevant characteristics from annotators and gather corresponding population-level data.
-National censuses and large-scale surveys could provide useful population benchmarks.
-[^4] Collection and release of annotator characteristics or weights derived from them may raise concerns about annotator confidentiality.
-This topic is outside the scope of this paper, however, the survey literature contains useful approaches (see Karr, 2016, for a review).
-We also note that collecting annotator characteristics may require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al.
-(2024).
-3) Implement data adjustment methods that account for differences between annotator and population characteristics.
-While our simple duplication approach showed promise, more sophisticated statistical techniques from survey research may yield better results.
+  3. 3)
+    Calculate weights that match the annotators to the population on those characteristics (Bethlehem et al., 2011; Valliant et al., 2013).
+  4. 4)
+    Use these weights in model training. Our simple duplication approach showed promise, future work should test more sophisticated weighting approaches.
 Limitations
-Stylized Biases.
+Stylized Biases and Simulated Data.
 Our simulation makes strong assumptions about annotator behavior, particularly in modeling consistent biases across annotator types.
@@
 We have not done that in this preliminary study because of the high cost and time needed to fine tune many RoBERTa models.
-As discussed, we have used five seeds in model training.
+As discussed, we used five seeds in model training.
+Need for Population Benchmarks and Annotator Characteristics.
+PAIR requires high quality benchmark information about the relevant population.
+These benchmarks might come from national statistical offices or national surveys.
+Annotators must provide accurate data on the same characteristics available in the benchmark data.
+In addition, theory from survey science demonstrates that bias will be reduced only when the characteristics used in weighting correlate with the annotations (Eckman et al., 2024).
+In our simulation, differences in annotations were driven solely by group membership ($A$, $B$).
+In the real world, it is challenging to know what characteristics impact annotation behavior and to find good benchmarks for those characteristics.
 Generalization Beyond Task Types.
@@
 Evaluation Metrics.
-While we measured calibration and classification accuracy, we did not examine other important metrics such as fairness across subgroups or robustness to adversarial examples.
-Future work should on training data adjustment should assess a broader range of performance measures.
+While we measured calibration and accuracy, we did not examine other important metrics such as fairness across subgroups or robustness to adversarial examples.
+Future work on training data adjustment should assess a broader range of performance measures.
 Ethical Considerations
@@
 BP is supported by ERC Consolidator Grant DIALECT (101043235).
-Appendix A Appendix
-A.1 Data
-Figure 5 shows the percentage of tweets labeled OL and HS in the four datasets for each value of $\beta$.
-The percentage in the Adjusted dataset is similar to that in the Balanced dataset for all values of $\beta$.
-The two unbalanced datasets have lower rates of OL and HS, because they overrepresent the A annotators, who are less likely to label OL and HS.
-Figure 5: Percentage of OL, HS “yes” labels, by dataset and bias ($\beta$)
-Because HS is rare in our dataset, as $\beta$ increases, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ values increase.
-This issue leads to a higher proportion of “yes” HS labels in the Balanced and Adjusted datasets, which have more B labels than the unadjusted datasets.
-[^5] If we had picked the B labels to overrepresent, we would have seen a converse problem with the OL labels.
-A.2 Model Training Details
+Appendix A PAIR Algorithm Implementation Details
+The adjustment we use to make the annotator pool more representative of the target population is a form of pseudo-population generation (Quatember, 2015).
+We create the pseudo-population by first constructing post-stratification weights, performing weight normalization to ensure the sum of the weights equals the size of the target population, and then duplicating each observation proportionally to its weight via deterministic replication.
+Post-stratification is a method of statistical adjustment that makes a selected sample more closely resemble a target population (Bethlehem et al., 2011; Valliant et al., 2013).
+We form strata (groups) of annotators.
+For example, we might form 6 strata from three categories of region and two categories of age.
+Post-stratification requires population-level totals or proportions for each stratum and corresponding case-level observations.
+The weight for each unit $i$ in stratum $s$ is:
+$w_{i}=\frac{P_{s}}{S_{s}}$ (4)
+where $P_{s}$ is the true population proportion (or total) for stratum $s$ and $S_{s}$ is the sample proportion (or total) for stratum $s$.
+In our case, the strata of interest was a single variable (annotator Types A & B).
+However, post-stratification can involve multiple variables if their joint distribution is known at the population level.
+Although the post-stratified weights will preserve the ratios of the strata in the target population, the weighted totals themselves may not match those in the target population.
+Weight normalization can be used to address this by updating the survey weights so that they sum to a desired total.
+The normalized weight for unit $i$ can be calculated by:
+$w_{i}^{\text{normalized}}=w_{i}^{\text{initial}}\cdot\frac{T}{\sum_{i=1}^{n}w_{i}^{\text{initial}}}$ (5)
+where $T$ is the target total. Since we want the Adjusted dataset to match the size of our gold standard Representative dataset, the target total for the simulation was 12 labels per tweet.
+Lastly, to construct a pseudo-population from our weighted data, we perform deterministic replication by replicating each unit $n_{a}$ times where:
+$n_{i}=\text{round(}w_{i}^{\text{normalized}}\text{)}$ (6)
+Rounding the normalized weights ensures that $n_{a}$ are integers and prevents fractional replicates.
+In this initial work, we prioritized this approach for its simple interpretation and reproducibility.
+However, researchers may prefer other approaches, such as replication via resampling, if they are interested in how the adjustment varies across samples, or incorporating weights directly into the loss function if they want to avoid the normalization and rounding steps.
+In our simulation, after post-stratification and weight normalization, each Type A label in the Non-representative 1 dataset receives a weight of 1 and each Type B label receives a weight of 2.
+This resulted in an Adjusted dataset where each of the Type A labels stays the same and each of the Type B labels is duplicated once.
+Appendix B Model Training Details
 Our implementation of RoBERTa models was based on the libraries pytorch (Paszke et al., 2019) and transformers (Wolf et al., 2020).
@@
 Table 3: Hyperparameter settings of RoBERTa models
-A.3 Additional Results
-We present the results of F1 scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$) in Figure 6.
-Figure 6: Model F1 scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$), by dataset and bias ($\beta$)
-A.4 Additional Metrics
-We also use the Expected Calibration Error (ECE) metric to evaluate model calibration.
-This metric quantifies calibration quality by measuring the weighted average absolute difference between accuracy and predicted confidence across $M$ bins:
-$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\cdot\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (4)
-where $B_{m}$ is the set of samples in the $m$-th bin, $|B_{m}|$ its size, $n$ the total number of samples, $\text{acc}(B_{m})$ the accuracy, and $\text{conf}(B_{m})$ the mean predicted confidence in $B_{m}$.
-Figure 7 shows the results.
-Consistent with our main findings in §4, we observe that for OL labels, the ECE values are relatively stable on the Balanced and Adjusted datasets ($t=-0.48,p=0.64$), with divergence with $\beta=0.15$.
-This contrasts with the significant divergence in ECE values for the Unbalanced subsets as the $\beta$ value increases.
-Figure 7: Model ECE metrics, by dataset and bias ($\beta$)
-Additionally, for model performance, we use another metric ROC-AUC in Figure 8. The results are very similar to the F1 results in §4.
-Figure 8: Model ROC-AUC metrics, by dataset and bias ($\beta$)
-A.5 Adjustment Details
-The adjustment we use to make the labeler pool more representative of the target population is a form of pseudo-population generation (Quatember, 2015).
-We create the pseudo-population by first constructing post-stratification weights, performing weight normalization to ensure the sum of the weights equals the size of the target population, and then duplicating each observation proportionally to its weight via deterministic replication.
-Post-stratification Bethlehem et al.
-(2011b) is a method of statistical adjustment that makes a selected sample more closely resemble a target population.
-Post-stratification requires population level totals or proportions and corresponding case-level observations of the same characteristic.
-to be available for each sample stratum that will be used in the weighting.
-The weight for unit $i$ is determined for each stratum by:
-$w_{i}=\frac{P_{s}}{S_{s}}$ (5)
-where $P_{s}$ is the true population proportion (or total) for stratum (or group) $s$ and $S_{s}$ is the sample proportion (or total) for stratum $s$.
-In our case, the strata of interest was a single variable (labeler Types A & B).
-However, post-stratification can involve multiple variables if their joint distribution is known at the population level.
-Although the post-stratified weights will preserve the ratios of the strata in the target population, the weighted totals themselves may not match those in the target population.
-Weight normalization can be used to address this by updating the survey weights so that they sum to a desired total.
-The normalized weight for unit $i$ can be calculated by:
-$w_{i}^{\text{normalized}}=w_{i}^{\text{initial}}\cdot\frac{T}{\sum_{i=1}^{n}w_{i}^{\text{initial}}}$ (6)
-where $T$ is the target total. Since we want the Adjusted dataset to match the size of our gold standard Balanced dataset, the target total for the simulation was 12 labels per tweet.
-Lastly, to construct a pseudo-population from our weighted data, we perform deterministic replication by replicating each unit $n_{i}$ times.
-In this initial work, we appreciated the simple interpretation and reproducibility of this approach.
-However, researcher may prefer other approaches, such as replication via resampling, if they are interested in how the adjustment varies across samples.
-In our case, after post-stratification and weight normalization, each Type A label in the Unbalanced 1 dataset receives a weight of 1 and each Type B label receives a weight of 2.
-This resulted in an Adjusted dataset where each of the Type A labels stays the same and each of the Type B labels is duplicated once.
+Appendix C Ablation Study 1: HS Results
+We show the results on HS labels in this section.
+We construct the HS probabilities for the A and B annotators are defined in the same way as the OL probabilities: $p_{i,HS}^{A}=\text{max}(p_{i,HS}-\beta,0)$, $p_{i,HS}^{B}=\text{min}(p_{i,HS}+\beta,1)$.
+We also construct the four data sets (Representative, Non-representative 1, Non-representative 2, Adjusted) in the same way we did for the OL case.
+Label Distribution during Simulations.
+Figure 4 shows the percentage of tweets labeled HS in the four datasets for each value of $\beta$.
+As expected, the percentage in the Adjusted dataset is similar to that in the Representative dataset for all values of $\beta$, and the two unbalanced datasets have lower rates of HS, because they overrepresent the A annotators, who are less likely to label HS.
+Figure 4: Percentage of tweets annotated as HS, by dataset and bias ($\beta$)
+HS is rare in our dataset (16.7% of HS labels are yes, Table 1).
+And, our simulation strategy overrepresents $A$ annotators in the two Non-representative data sets, who are less likely to perceive HS (Table 2).
+For these reasons, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ probabilities continue to increase.
+This issue leads the proportion of “yes” labels in the Representative and Adjusted datasets to increase with $\beta$ in the HS data set, which have more B labels than the unadjusted datasets.
+Results.
+Figure 5 (1) contains the ACB results and Figure 5 (2) the F1 score results for the HS datasets.
+The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models.
+This effect is likely due to the combination of label rarity and our simulation design.
+With few positive labels to begin with, the impact of varying annotator characteristics through our $\beta$ parameter may be overwhelmed by the baseline scarcity of hate speech annotations.
+Ablation Study 2 (Appendix §D) addresses this point.
+Figure 5: Metrics for HS models, by dataset and bias ($\beta$)
+Appendix D Ablation Study 2: Difficult Tweets
+Our simulations assumed that all tweets are impacted the same way (Eq.
+(2)), which is an oversimplification.
+More realistically, annotator characteristics likely have more impact for ambiguous tweets.
+For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$.
+This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq.
+2).
+The filtered datasets contain 267 (OL) and 360 (HS) tweets.
+Figures 6 and 7 show the results for two metrics (ACB, F1) for filtered OL and HS labels.
+In Figure 6(1): the Representative and Adjusted models have similar ACB and are lower than the Non-representative models.
+The F1 scores do not show differences between the models.
+These results are similar to those on the full set of tweets (Figure 3).
+In the two HS figures (7), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics.
+These results are more promising than those on the full set of tweets (Figure 5).
+Figure 6: Metrics for OL models for filtered tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
+Figure 7: Metrics for HS models for filtered tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)

### v3

@@
-Correcting Annotator Bias in Training Data: Population-Aligned Instance Replication (PAIR)
+Aligning NLP Models with Target Population Perspectives using PAIR: Population-Aligned Instance Replication
 Abstract
-Models trained on crowdsourced labels may not reflect broader population views, because those who work as annotators do not represent the population.
-We propose Population-Aligned Instance Replication (PAIR), a method to address bias caused by non-representative annotator pools.
-Using a simulation study of offensive language and hate speech, we create two types of annotators with different labeling tendencies and generate datasets with varying proportions of the types.
-We observe that models trained on unbalanced annotator pools show poor calibration compared to those trained on representative data.
-By duplicating labels from underrepresented annotator groups to match population proportions, PAIR reduces bias without collecting additional annotations.
-These results suggest that statistical techniques from survey research can improve model performance.
-We conclude with practical recommendations for improving the representativity of training data and model performance.
+Models trained on crowdsourced annotations may not reflect population views, if those who work as annotators do not represent the broader population.
+In this paper, we propose PAIR: Population-Aligned Instance Replication, a post-processing method that adjusts training data to better reflect target population characteristics without collecting additional annotations.
+Using simulation studies on offensive language and hate speech detection with varying annotator compositions, we show that non-representative pools degrade model calibration while leaving accuracy largely unchanged.
+PAIR corrects these calibration problems by replicating annotations from underrepresented annotator groups to match population proportions.
+We conclude with recommendations for improving the representativity of training data and model performance.
 [^1] The code for experiments is available at https://github.com/soda-lmu/PAIR.
-Correcting Annotator Bias in Training Data:
-Population-Aligned Instance Replication (PAIR)
+Aligning NLP Models with Target Population Perspectives using PAIR:
+Population-Aligned Instance Replication
 […]
 1 Introduction and Inspiration
-When a hate speech detection model flags harmless expressions as toxic, or a content moderation system fails to identify genuinely harmful content, the root cause often lies not in the model architecture, but in who labeled the training data.
-While NLP models aim to serve broad populations, the human judgments used to train these systems typically come from a non-representative pool of annotators – crowdworkers and convenience samples whose demographics, cultural contexts, and worldviews may differ from the communities the models ultimately impact Sorensen et al.
+When a hate speech detection model flags harmless expressions as toxic, or a content moderation system fails to identify genuinely harmful content, the root cause often lies not in the model architecture, but in who annotated the training data.
+While Natural Language Processing (NLP) models aim to serve broad populations, the human judgments used to train these systems often come from crowdworkers and convenience samples.
+And the demographics, cultural contexts, and worldviews of these annotators often differ from those of the communities the models ultimately impact Sorensen et al.
 (2024); Fleisig et al.
 (2024).
-These non-representative annotator pools can have real consequences, because annotator characteristics like age, education level, and cultural background impact how content is labeled Sap et al.
+These non-representative annotator pools can have real consequences, because annotator characteristics like age, education level, and cultural background impact how content is annotated Sap et al.
 (2022); Fleisig et al.
@@
 (2024).
-Models trained on non-representative data can perpetuate the biases and blind spots of their training data Smart et al.
-(2024); Berinsky et al.
-(2012); Ouyang et al.
+The influence of annotator characteristics underscores that language understanding is not a single objective truth but a constellation of equally valid interpretations anchored in different lived experiences.
+When this perspectivist interpretation is ignored, models trained on non-representative data can perpetuate the biases and blind spots of their limited training data Berinsky et al.
+(2012); Hebert-Johnson et al.
 (2018); Mehrabi et al.
 (2021); Rolf et al.
+(2021); Hüllermeier and Waegeman (2021); Ouyang et al.
 (2022); Favier et al.
-(2023); Hebert-Johnson et al.
-(2018); Hüllermeier and Waegeman (2021).
-Figure 1: Top: Adjusting survey data to match population produces high quality results. Bottom: Can a similar adjustment in data annotations also improve model performance?
-Fortunately, survey researchers have developed robust statistical techniques to produce population-level estimates from non-representative samples Bethlehem et al.
+(2023); Smart et al.
+(2024).
+Figure 1: Top: Adjusting survey data to match population produces high quality results.
+Bottom: Can a similar adjustment in data annotations also improve model performance?
+Fortunately, survey researchers have developed statistical techniques to produce population-level estimates from non-representative samples Bethlehem et al.
 (2011).
 The top panel of Figure 1 shows a simple survey workflow: collecting survey data, creating statistical weights to match the sample to the population, and estimating population parameters.
-We propose adapting this approach to the machine learning context, enabling models to better align with target populations even when trained on non-representative annotator pools (bottom panel).
-To test this approach, we conduct a rigorous simulation study following established practices in statistical research (Burton et al., 2006; Valliant, 2019; Morris et al., 2019).
-Using controlled experiments, simulated populations, and multiple training datasets varying in annotator composition, we investigate two questions:
+We adapt this approach to the machine learning context, enabling models to better align with target populations even when trained on non-representative annotator pools (bottom panel).
+Our Population-Aligned Instance Replication (PAIR) method post-processes training data to better reflect target populations without collecting additional annotations.
+We test the approach with a simulation study (Burton et al., 2006; Valliant, 2019; Morris et al., 2019) and answer two questions:
   • RQ1: How do non-representative annotator pools impact model calibration and accuracy?
-  • RQ2: Can survey techniques effectively mitigate these annotator pool effects?
-Our results demonstrate that models trained on nonrepresentative annotator pools perform poorly.
+  • RQ2: Can our proposed weighting method (PAIR) mitigate these annotator pool effects?
+Our results demonstrate that models trained on non-representative annotator pools perform worse than those trained on representative pools.
 However, simple adjustment methods can improve performance without collecting additional data.
-These findings suggest that insights from survey methodology can help AI systems better represent the populations they serve.
+These findings suggest that insights from survey methodology can help artificial intelligence (AI) systems better represent the populations they serve.
 2 Related Work
@@
 Annotator Impact on Data and Models.
-Annotator characteristics and attitudes significantly influence label quality, particularly for subjective tasks like toxicity detection (Giorgi et al., 2024; Prabhakaran et al., 2021; Fleisig et al., 2023; Sap et al., 2022).
+Annotator characteristics and attitudes significantly influence annotation quality, particularly for subjective tasks like toxicity detection (Giorgi et al., 2025; Prabhakaran et al., 2021; Fleisig et al., 2023; Sap et al., 2022).
 For example, annotators’ political views and racial attitudes affect their toxicity judgments (Sap et al., 2022).
-Models trained on non-representative annotator pools inherit these biases and generalize poorly (Smart et al., 2024; Berinsky et al., 2012; Ouyang et al., 2022; Mehrabi et al., 2021; Rolf et al., 2021; Favier et al., 2023).
+Models trained on non-representative annotator pools inherit these biases and generalize poorly (Berinsky et al., 2012; Mehrabi et al., 2021; Rolf et al., 2021; Ouyang et al., 2022; Favier et al., 2023; Smart et al., 2024; Mokhberian et al., 2024).
 Annotator Demographics.
-Several researchers advocate collecting annotator demographics to assess representation and identify biases (Bender and Friedman, 2018; Prabhakaran et al., 2021).
-[^2] In our context, these characteristics are used only to analyze bias; because they are not available for unlabeled text, they are not features that the model can use.
+Several researchers advocate collecting annotator demographics to assess representation and identify biases (Bender and Friedman, 2018; Prabhakaran et al., 2021; Plank, 2022; Wan et al., 2023; Santy et al., 2023; Pei and Jurgens, 2023).
+[^2] In our context, these characteristics are used only to analyze bias.
+Because they are not available for unannotated text, they are not features that the model can use.
 However, collecting and releasing these data can raise privacy concerns (Fleisig et al., 2023).
+Recent works have also used demographics to prompt the large language models Argyle et al.
+(2023), and some find that these are less effective in subjective contexts Sun et al.
+(2025); Orlikowski et al.
+(2025).
 Debiasing & Data Augmentation Methods.
-Prior work has proposed various approaches to reduce bias in training data features and labels (Calders et al., 2009; Kamiran and Calders, 2012; Sharma et al., 2020).
+Prior work has proposed various approaches to reduce bias in training data features and annotations.
 Most similar to our work is the resampling and reweighting approaches of Calders et al.
-(2009) and Kamiran and Calders (2012), and the oversampling of minority class cases of Ling and Li (1998).
+(2009) and Kamiran and Calders (2012), imputation Lowmanstone et al.
+(2023), and the oversampling of minority class cases of Ling and Li (1998).
 PAIR adapts these methods to balance annotator characteristics rather than class labels or sensitive observation-level features.
-PAIR retains the simplicity and interpretability of earlier resampling methods while extending them to a “Learning with Disagreement” (Uma et al., 2021; Leonardelli et al., 2023) setting with multiple annotations per observation, by replicating labels from underrepresented annotator groups.
+PAIR retains the simplicity and interpretability of earlier resampling methods while extending them to a “Learning with Disagreement” (Uma et al., 2021; Leonardelli et al., 2023) setting with multiple annotations per observation, by replicating annotations from underrepresented annotator groups.
+3 PAIR Algorithm: Adjustment via Pseudo-Population
+To adjust an annotator pool to better reflect a target population, we propose the PAIR algorithm, which constructs a pseudo-population through post-stratification, weight normalization, and deterministic replication.
+This adjustment strategy is inspired by established methods in survey sampling (Quatember, 2015).
+Post-stratification aligns a sample more closely with population-level distributions (Bethlehem et al., 2011; Valliant et al., 2013).
+Annotators are grouped into strata based on demographic or behavioral characteristics.
+For each unit $i$ in stratum $s$, a post-stratification weight is computed as:
+$w_{s,i}=\frac{P_{s}}{S_{s}}$ (1)
+where $P_{s}$ and $S_{s}$ denote the share of the population in stratum $s$ and the share of the sample (or annotator pool), respectively.
+The $P$ values come from official statistics or surveys.
+The $S$ values likely come from the annotators themselves and researchers may have to collect them.
+This technique can accommodate multiple stratification variables; it is only limited by the availability of population or reference data and data about the annotators.
+These weights have only relative meaning and are invariant to multiplication by a constant ($K$):
+$w_{i}^{\text{normalized}}=w_{i}^{\text{initial}}\times K$ (2)
+Normalization useful if research teams have a target number of annotations per observation in mind, for either computational or design reasons, or if some weights given by Eq.
+1 are very small and round to one.
+To generate a pseudo-population, we apply deterministic replication: each unit is replicated $n_{i}$ times where
+$n_{i}=\text{round}(w_{i}^{\text{normalized}})-1$ (3)
+ensuring integer replication counts. This approach produces a dataset that reflects population proportions while maintaining interpretability and reproducibility.
+While we focus in this initial study on deterministic replication, alternative implementations are possible, including resampling-based replication or direct incorporation of weights into model training.
 4 Annotation Simulation and Model Training
-To address our research questions, we imagine a population made up of two types of people: those more likely to perceive offensive language and hate speech and those less likely.
-We create three sets of simulated annotations which differ in the mix of the annotator types.
-We then create a fourth data set, using the PAIR algorithm, to attempt to fix the imbalance in the annotators.
-We fine-tune RoBERTa models on the four data sets and evaluate the effect of annotator characteristics on model performance (RQ1) and the ability of the PAIR algorithm to improve performance (RQ2).
+To address our research questions, we conduct a simulation study on offensive language and hate speech detection.
+We imagine a population made up of equal shares of two types of people: those more likely to perceive offensive language and hate speech and those less likely.
+We create three datasets of simulated annotations which differ in the mix of the annotator types.
+We then create a fourth dataset, using the PAIR algorithm, to fix the imbalance in the annotators.
+We fine-tune RoBERTa models on the four datasets and evaluate the effect of annotator composition on model performance (RQ1) and the ability of the PAIR algorithm to improve performance (RQ2).
 4.1 Simulating Annotations
-We use a dataset of 3,000 English-language tweets, each with 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no) (Kern et al., 2023).
-[^3] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2 We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with.
-We randomly select (without replacement) 12 labels (of both OL and HS) of each tweet in the original dataset.
-[^4] As shown in Table 2, we can more carefully control the construction of our data sets when the number of labels per tweet is divisible by four.
-Let $p_{i,OL}$ be the proportion of the 12 annotators who labeled tweet $i$ as OL and $p_{i,HS}$ defined similarly.
-Table 1 shows the distribution of these proportions across the 3,000 tweets.
-The HS labels are unevenly distributed, whereas the OL labels are relatively balanced.
-Because this work is a preliminary investigation of the PAIR approach, and balanced labels provide a clearer view of PAIR’s abilities, the main body of the paper focuses on results with the OL labels; the HS results in Appendix §C.
-Variable $25^{th}$ pctile. Median Mean $75^{th}$ pctile.
-$p_{i,OL}$ 0.167 0.667 0.564 0.917
-$p_{i,HS}$ 0.083 0.167 0.301 0.50
-Table 1: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in original data
-We imagine a population made up of equal shares of two types of people. Type A people are less likely to say a tweet contains OL. Type B people are more likely:
+We use our previously collected dataset on tweet annotation sensitivity (Kern et al., 2023) [^3] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2, which is a dataset of 3,000 English-language tweets, each with 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no).
+We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with.
+We randomly select (without replacement) 12 annotations (of both OL and HS) of each tweet in the original dataset.
+[^4] As shown in Table 1, we can more carefully control the construction of our datasets when the number of annotations per tweet is even.
+Let $p_{i,OL}$ be the proportion of the 12 annotators who annotated tweet $i$ as OL and $p_{i,HS}$ defined similarly.
+Figure 2 shows the distribution of these proportions across the 3,000 tweets.
+The HS annotations are clustered near 0, whereas the OL annotations are more spread out between 0 and $1$.
+Figure 2: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in original data
+The population contains two types of people (50% each). Type A people are less likely to say a tweet contains OL. Type B people are more likely:
 $p_{i,OL}^{A}$ $=\text{max}(p_{i,OL}-\beta,0)$ (4)
@@
 This range is large on the probability scale and covers most reasonable situations.
-With these seven values of $\beta$, we create seven vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B}$) for each tweet.
+With these six values of $\beta$, we create six vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B}$) for each tweet.
 We then create four datasets, each with 3,000 tweets (Table 1), for each value of $\beta$.
-The Representative Dataset contains OL labels from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$).
-Because the proportion of A and B annotators in this dataset matches the population, the labels in this dataset are our gold standard.
-Dataset Labels per tweet A labels B labels
+The Representative Dataset contains OL annotations from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$).
+The proportion of A and B annotators in this dataset matches the simulated population we created.
+Annotations A B
+Dataset per tweet annotations annotations
 Representative 12 6 6
@@
 Adjusted 12 6 3 + 3*
-* 3 B labels duplicated
+* 3 B annotations replicated
 Table 1: Four training datasets for each bias value ($\beta$)
-We then create two unbalanced datasets.
-Non-representative 1 randomly deletes three B labels for each tweet from the Representative Dataset.
-Non-representative 2 adds three additional A labels, drawn from $p_{i}^{A}$, to the Non-representative 1 dataset.
+We next create two unbalanced datasets.
+Non-representative 1 randomly deletes three B annotations for each tweet from the Representative Dataset.
+Non-representative 2 adds three additional A annotations, drawn from $p_{i,OL}^{A}$, to the Non-representative 1 dataset.
 The Non-representative 2 Dataset is more unbalanced than Non-representative 1, but contains the same number of annotations as the Representative dataset.
+4.2 Applying PAIR Algorithm
 Finally, we use the PAIR algorithm to create the Adjusted Dataset.
-It is the same as the Non-representative 1 dataset, but the three B annotations are duplicated.
-This duplication is an easy way to adjust the unbalanced training dataset to reflect the population.
-Appendix §A provides a general version of the PAIR algorithm which can handle imbalances across multiple annotator characteristics.
-Figure 2 shows the percentage of tweets labeled OL in the four datasets for each value of $\beta$.
-Figure 2: Percentage of tweets annotated as OL, by dataset and bias ($\beta$)
+Starting with the Non-representative 1 Dataset, we calculate the share of the annotator pool that is in the A and B strata: $S_{A}=\frac{2}{3},S_{B}=\frac{1}{3}$.
+The population proportions, by construction, are $P_{A}=0.5,P_{B}=0.5$.
+Applying (1), we get $w_{A,i}=0.75,w_{B,i}=1.5$.
+We multiply these weights by $K=\frac{4}{3}$ to get $w_{A,i}=1,w_{B,i}=2$.
+These weights give us $n_{A,i}=0,n_{B,i}=1$, which leads us to replicate all B annotations in the Non-representative 1 Dataset (see Table 1).
+The HS probabilities for the A and B annotators are defined in the same way: $p_{i,HS}^{A}=\text{max}(p_{i,HS}-\beta,0)$, $p_{i,HS}^{B}=\text{min}(p_{i,HS}+\beta,1)$.
+We also construct the four datasets (Representative, Non-representative 1, Non-representative 2, Adjusted) in the same way we did in the OL case.
+Figures 3 and 4 show the percentage of instances annotated OL and HS in the four datasets for each value of $\beta$.
+In both, the percentage of OL/HS annotations in the Adjusted dataset is similar to that in the Representative dataset for all values of $\beta$.
+The percentage in the two unbalanced datasets is lower, because those datasets overrepresent the A annotators, who are less likely to annotate OL/HS.
+HS is rare in our dataset (16.7% of instances were annotated as HS), and our simulation strategy overrepresents A annotators in the two Non-representative datasets, who are less likely to perceive HS (Table 1).
+For these reasons, as $\beta$ increases, more $p_{i,HS}^{A}$ are 0 while the $p_{i,HS}^{B}$ probabilities increase.
+This issue leads the proportion of HS annotations in the Representative and Adjusted datasets to increase with $\beta$ in the HS dataset, which have more B annotations than the unadjusted datasets (Figure 4).
+Figure 3: Percentage of instances annotated as OL, by dataset and bias ($\beta$)
+Figure 4: Percentage of instances annotated as HS, by dataset and bias ($\beta$)
 4.3 Model Training and Evaluation
 Training and Test Setup.
-We train models on each dataset.
-We divide each dataset, at the tweet level, into training (2000 tweets), development (500), and test (500) sets.
+We train models on each of the eight datasets: four for OL, four for HS.
+We divide each dataset, at the tweet level, into training (2,000 tweets), development (500), and test (500) sets.
 Each tweet appears 12 times in the Representative, Non-representative 2, and Adjusted datasets and nine times in the Non-representative 1 set.
@@
 To ensure reliable results, we trained five versions with different random seeds and averaged their performance.
-Appendix §B contains details on model training.
+Our implementation of RoBERTa models was based on the libraries pytorch (Paszke et al., 2019) and transformers (Wolf et al., 2020).
+During training, we used the same hyperparameter settings (Table 2) for the five training conditions to keep these variables consistent for comparison purposes.
+We trained each model variation with five random seeds $\{10,42,512,1010,3344\}$ and took the average across the models.
+All experiments were conducted on an NVIDIA® A100 80 GB RAM GPU.
+Hyperparameter Value
+encoder roberta-base
+epochs_trained 55
+learning_rate $3e^{-5}$
+batch_size $32$
+warmup_steps $500$
+optimizer AdamW
+max_length $128$
+Table 2: Hyperparameter settings of RoBERTa models
 Performance Metrics.
-We evaluate models using both calibration and accuracy metrics on the test set.
+We evaluate models using calibration and accuracy metrics on the test set.
 While accuracy metrics directly measure classification performance, calibration metrics provide crucial insights into model reliability by assessing probability estimate quality – particularly important for high-stakes applications requiring trustworthy confidence measures.
-For calibration, we report Absolute Calibration Bias (ACB), which measures how well model predicted probabilities match true frequencies in the Representative dataset (lower is better).
-$\text{ACB}_{OL}=\frac{1}{n}\sum_{i=1}^{n}\left|p_{i,OL}-\text{preds}_{i,OL}\right|$ (3)
+For calibration, we report Absolute Calibration Bias (ACB, Equation 6), which measures how well a model’s predicted probabilities align with true annotation frequencies.
+For each tweet $i$, we compare the model’s predicted probability of offensive language ($\text{preds}_{i,OL}$) against the true proportion of annotators who labeled that tweet as offensive ($p_{i,OL}$).
+$\text{ACB}_{OL}=\frac{1}{n}\sum_{i=1}^{n}\left|\text{preds}_{i,OL}-p_{i,OL}\right|$ (6)
+$\text{ACB}_{HS}$ is defined accordingly.
+ACB adapts established calibration metrics by using the annotator agreement proportion as a plug-in estimator for the true probability, avoiding the need for binning (as in ECE, Naeini et al., 2015) while maintaining the intuitive L1 distance interpretation (Roelofs et al., 2022).
+A low ACB score indicates that the model’s confidence scores accurately reflect the underlying annotation uncertainty in the population.
 For accuracy, we report the F1 score.
 5 Results
+We show results separately for the OL and HS models.
+5.1 OL Models
 Calibration.
-Figure 3(1) compares the ACB in the test set for models trained on the simulated datasets. The dark lines show average ACB across the five training runs and the shading shows the standard deviation.
-The ACB for the models trained on the Adjusted dataset closely tracks that for the Representative (gold standard) data set and does not increase with $\beta$.
+Figure 5 compares the ACB in the test set for models trained on the four datasets. The dark lines show average ACB across the five training runs and the shading shows the standard deviation.
+The ACB for the models trained on the Adjusted dataset closely tracks that for the Representative dataset and does not increase with $\beta$.
 ACB for the models trained on the two unbalanced datasets is greater and grows with $\beta$.
 These results demonstrate the effectiveness of our adjustment method.
-Duplicating the labels from the underrepresented annotator type to match population proportions improves model calibration.
-Figure 3: Metrics for OL models, by dataset and bias ($\beta$)
+Replicating the annotations from the underrepresented annotator type to match population proportions improves model calibration.
+Figure 5: ACB scores for OL Models, by dataset and bias ($\beta$)
+Figure 6: F1 scores for OL Models, by dataset and bias ($\beta$)
+Figure 7: ACB scores for HS Models, by dataset and bias ($\beta$)
+Figure 8: F1 scores for HS Models, by dataset and bias ($\beta$)
+Figure 9: Percentage of OL instances on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$) by dataset and bias ($\beta$)
+Figure 10: Percentage of HS instances on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$) by dataset and bias ($\beta$)
 Accuracy.
-Figure 3(2) compares the models’ F1 scores.
-In contrast to Figure 3(1), we do not see strong differences between the models trained on the different datasets.
-For all datasets, model performance declines with $\beta$: as the amount of bias in the labels increases, the models have a harder time predicting the binary OL label.
+Figure 6 compares the models’ F1 scores.
+In contrast to Figure 5, we do not see strong differences between the models trained on the different datasets.
+For all datasets, model performance declines with $\beta$: as the amount of bias in the annotations increases, the models are less able to predict the binary OL label.
 Because the F1 metric focuses on binary predictions, it is less sensitive to training biases than calibration metrics like ACB, which more explicitly capture biases through prediction scores.
-These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models: binary classification metrics can obscure such effects.
 In decision-making, miscalibrated predictions can have harmful consequences when, for example, hateful content remains undetected Van Calster et al.
 (2019).
-Ablation Study 1: Hate Speech labels.
-Appendix §C gives the ACB and F1 results for the HS labels.
-The PAIR approach has limited effectiveness when label distributions are highly skewed.
-This failure is likely due to the rarity of hate speech labels in the dataset (only 16.7% positive labels) combined with our simulation setup, which further reduces positive labels in the unbalanced datasets.
-PAIR may need modifications, such as stratified sampling or adaptive replication, when working with highly imbalanced datasets.
-Ablation Study 2: Difficult Tweets.
-A representative pool of annotators may be more important for instances that are difficult or about which people disagree.
-Filtering our tweets in this way also removes cases where floor and ceiling effects could mask the impact of annotator characteristics.
-To test PAIR performance on difficult tweets, we reran the models on tweets where $0.60.4<p_{i,HS}<$ 0.6, $0.60.4<p_{i,OL}<$ 0.6.
-Performance on OL is still good, and performance on HS is improved (Appendix §D).
+These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models: binary classification metrics can obscure such effects.
+5.2 HS Models
+Figure 7 contains the ACB results and Figure 8 the F1 score results for the HS models trained on each dataset.
+Though the adjusted model roughly tracks the representative models for ACB, there is instability in the results.
+All models show lower average ACB values than the representative model across a wide range of the bias offset (0.10 - 0.20).
+The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models.
+This effect is likely due to the combination of label rarity and our simulation design.
+With few positive annotations to begin with, the impact of the $\beta$ parameter and the overrepresentation of the A annotators may be overwhelmed by the baseline scarcity of hate speech annotations.
+Calibration metrics can be less reliable with rare classes Zhong et al.
+(2021).
+5.3 Sensitivity Analysis: Difficult Tweets
+Our simulations assumed that all annotator type impacts all tweets the same way (Eq.
+5), which is an oversimplification.
+More likely, annotator characteristics have more impact for ambiguous tweets.
+For example, prior research in the psychology literature on judgment under uncertainty suggests that people draw more heavily on personal heuristics when interpreting unclear or underspecified information Tversky and Kahneman (1974).
+For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$.
+Subsetting the tweets in this way also eliminates the floor and ceiling effects in Eq.
+5.
+The filtered datasets contain 267 (OL) and 360 (HS) tweets.
+The proportions of OL and HS annotations are stable for the Representative and Adjusted sets and decrease for the Non-representative sets as we increase the bias offset (Figures 9 and 10).
+This mimics the trend for OL in the full dataset (Figure 3).
+Figure 11: ACB scores for OL Models, on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
+Figure 12: F1 scores for OL Models, on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
+Figures 11, 12, 13, and 14 show results for two metrics (ACB, F1) for filtered OL and HS annotations.
+In Figure 11, the Representative and Adjusted models have similar ACB and are lower than the Non-representative models.
+The F1 scores do not show differences between the models.
+These results are similar to those on the full set of tweets (Figures 5 and 6).
+In the two HS figures (13, 14), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics.
+These results are more promising than those on the full set of tweets (Figures 7 and 8) and support our hypothesis that the rarity of HS annotations contributed to the lack of positive results for the PAIR approach in §5.2.
+The PAIR algorithm works well with difficult tweets, which is where it is likely most needed.
+Figure 13: ACB scores for HS Models, on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)
+Figure 14: F1 scores for HS Models, on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)
 6 Discussion & Recommendations
-Our results show that (RQ1) OL prediction models perform less well when trained on data from non-representative annotator pools, and (RQ2) simple statistical adjustments can improve model calibration without collecting additional annotations.
-These findings establish a promising bridge between survey statistics and machine learning - offering a practical approach to make AI systems more representative of and responsive to the populations they serve, particularly for tasks involving subjective human judgments.
-We recommend these four steps to reduce bias due to non-representative annotator pools:
+Our experimental results show the OL prediction models perform less well when trained on data from non-representative annotator pools (RQ1), and simple statistical adjustments can improve model calibration without collecting additional annotations or involving additional annotators (RQ2).
+While PAIR’s impact was harder to assess for the rare HS class, PAIR did improve calibration of both the OL and HS models when trained on difficult tweets.
+These findings establish a promising bridge between survey statistics and machine learning – offering a practical approach to make AI systems more representative of and responsive to the populations they serve, particularly for tasks involving subjective human judgments.
+We recommend the following four steps to reduce bias due to non-representative annotator pools:
   1. 1)
@@
   2. 2)
-    Collect these characteristics from annotators and gathering corresponding population-level data from national censuses or high-quality surveys.
+    Collect these characteristics from annotators and gather corresponding population-level data from national censuses or high-quality surveys.
 [^5] Collection and release of annotator characteristics or weights derived from them may raise confidentiality concerns.
-The survey literature contains relevant approaches (see Karr, 2016, for a review).
+The survey literature offers advice for sharing sensitive data (see Karr, 2016, for a review).
 Collecting annotator characteristics may also require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al.
@@
   4. 4)
-    Use these weights in model training. Our simple duplication approach showed promise, future work should test more sophisticated weighting approaches.
-Limitations
+    Use these weights in model training. Our simple replication approach showed promise, and future work should test more sophisticated weighting approaches.
+7 Limitations
+Our study explores bias-aware data simulation and evaluation in a controlled setting, which necessarily involves simplifying assumptions and methodological constraints.
+We outline key areas where future work could broaden the applicability and robustness of our findings.
 Stylized Biases and Simulated Data.
-Our simulation makes strong assumptions about annotator behavior, particularly in modeling consistent biases across annotator types.
+Our simulation makes strong assumptions about annotator behavior: there are only two types of annotators, and, within each type, annotators behave similarly.
 Real-world annotator biases may be more nuanced or context-dependent.
+The simulated annotators might not be representative of a stable opinion group Mokhberian et al.
+(2024); Vitsakis et al.
+(2024).
 Future work could incorporate more realistic biases and refine the proposed simulations and statistical techniques.
 Sampling Variability.
-We have created only one version the four datasets for each label type and value of $\beta$, each of which contains random draws from the Bernoulli distribution.
+We have created only one version of the four datasets for each annotation type and value of $\beta$, each of which contains random draws from the Bernoulli distribution.
 A more traditional statistical approach would create multiple versions of the datasets and train models on each one, to average over the sampling variability.
-We have not done that in this preliminary study because of the high cost and time needed to fine tune many RoBERTa models.
+Though limited by computational constraints in this work, future work could take on a more expansive simulation.
 As discussed, we used five seeds in model training.
@@
 Annotators must provide accurate data on the same characteristics available in the benchmark data.
-In addition, theory from survey science demonstrates that bias will be reduced only when the characteristics used in weighting correlate with the annotations (Eckman et al., 2024).
+Unfortunately, annotators sometimes do not provide accurate information Chandler and Paolacci (2017); Huang et al.
+(2023).
+In addition, theory demonstrates that bias will be reduced only when the characteristics used in weighting correlate with the annotations (Eckman et al., 2024).
 In our simulation, differences in annotations were driven solely by group membership ($A$, $B$).
-In the real world, it is challenging to know what characteristics impact annotation behavior and to find good benchmarks for those characteristics.
+In the real world, it is challenging to know what characteristics impact annotation behavior for a given task and to find good benchmarks for those characteristics.
 Generalization Beyond Task Types.
@@
 Future work on training data adjustment should assess a broader range of performance measures.
-Ethical Considerations
-In this simulation study, we experiment on a publicly available dataset Kern et al.
+8 Ethical Considerations
+In this simulation study, we experiment on a publicly available dataset collected in our previous study Kern et al.
 (2023), which contains offensive and hateful tweets.
-We do not support the views expressed in the tweets.
+We do not support the views expressed in these tweets.
 The simulation study itself does not collect any new data or raise any ethical considerations.
@@
 BP is supported by ERC Consolidator Grant DIALECT (101043235).
-Appendix A PAIR Algorithm Implementation Details
-The adjustment we use to make the annotator pool more representative of the target population is a form of pseudo-population generation (Quatember, 2015).
-We create the pseudo-population by first constructing post-stratification weights, performing weight normalization to ensure the sum of the weights equals the size of the target population, and then duplicating each observation proportionally to its weight via deterministic replication.
-Post-stratification is a method of statistical adjustment that makes a selected sample more closely resemble a target population (Bethlehem et al., 2011; Valliant et al., 2013).
-We form strata (groups) of annotators.
-For example, we might form 6 strata from three categories of region and two categories of age.
-Post-stratification requires population-level totals or proportions for each stratum and corresponding case-level observations.
-The weight for each unit $i$ in stratum $s$ is:
-$w_{i}=\frac{P_{s}}{S_{s}}$ (4)
-where $P_{s}$ is the true population proportion (or total) for stratum $s$ and $S_{s}$ is the sample proportion (or total) for stratum $s$.
-In our case, the strata of interest was a single variable (annotator Types A & B).
-However, post-stratification can involve multiple variables if their joint distribution is known at the population level.
-Although the post-stratified weights will preserve the ratios of the strata in the target population, the weighted totals themselves may not match those in the target population.
-Weight normalization can be used to address this by updating the survey weights so that they sum to a desired total.
-The normalized weight for unit $i$ can be calculated by:
-$w_{i}^{\text{normalized}}=w_{i}^{\text{initial}}\cdot\frac{T}{\sum_{i=1}^{n}w_{i}^{\text{initial}}}$ (5)
-where $T$ is the target total. Since we want the Adjusted dataset to match the size of our gold standard Representative dataset, the target total for the simulation was 12 labels per tweet.
-Lastly, to construct a pseudo-population from our weighted data, we perform deterministic replication by replicating each unit $n_{a}$ times where:
-$n_{i}=\text{round(}w_{i}^{\text{normalized}}\text{)}$ (6)
-Rounding the normalized weights ensures that $n_{a}$ are integers and prevents fractional replicates.
-In this initial work, we prioritized this approach for its simple interpretation and reproducibility.
-However, researchers may prefer other approaches, such as replication via resampling, if they are interested in how the adjustment varies across samples, or incorporating weights directly into the loss function if they want to avoid the normalization and rounding steps.
-In our simulation, after post-stratification and weight normalization, each Type A label in the Non-representative 1 dataset receives a weight of 1 and each Type B label receives a weight of 2.
-This resulted in an Adjusted dataset where each of the Type A labels stays the same and each of the Type B labels is duplicated once.
-Appendix B Model Training Details
-Our implementation of RoBERTa models was based on the libraries pytorch (Paszke et al., 2019) and transformers (Wolf et al., 2020).
-During training, we used the same hyperparameter settings of the respective models for our five training conditions to keep these variables consistent for comparison purposes.
-We report the hyperparameter settings of the models in Table 3.
-To avoid random effects on training, we trained each model variation with five random seeds $\{10,42,512,1010,3344\}$ and took the average across the models.
-All experiments were conducted on an NVIDIA® A100 80 GB RAM GPU.
-Hyperparameter Value
-encoder roberta-base
-epochs_trained $5$
-learning_rate $3e^{-5}$
-batch_size $32$
-warmup_steps $500$
-optimizer AdamW
-max_length $128$
-Table 3: Hyperparameter settings of RoBERTa models
-Appendix C Ablation Study 1: HS Results
-We show the results on HS labels in this section.
-We construct the HS probabilities for the A and B annotators are defined in the same way as the OL probabilities: $p_{i,HS}^{A}=\text{max}(p_{i,HS}-\beta,0)$, $p_{i,HS}^{B}=\text{min}(p_{i,HS}+\beta,1)$.
-We also construct the four data sets (Representative, Non-representative 1, Non-representative 2, Adjusted) in the same way we did for the OL case.
-Label Distribution during Simulations.
-Figure 4 shows the percentage of tweets labeled HS in the four datasets for each value of $\beta$.
-As expected, the percentage in the Adjusted dataset is similar to that in the Representative dataset for all values of $\beta$, and the two unbalanced datasets have lower rates of HS, because they overrepresent the A annotators, who are less likely to label HS.
-Figure 4: Percentage of tweets annotated as HS, by dataset and bias ($\beta$)
-HS is rare in our dataset (16.7% of HS labels are yes, Table 1).
-And, our simulation strategy overrepresents $A$ annotators in the two Non-representative data sets, who are less likely to perceive HS (Table 2).
-For these reasons, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ probabilities continue to increase.
-This issue leads the proportion of “yes” labels in the Representative and Adjusted datasets to increase with $\beta$ in the HS data set, which have more B labels than the unadjusted datasets.
-Results.
-Figure 5 (1) contains the ACB results and Figure 5 (2) the F1 score results for the HS datasets.
-The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models.
-This effect is likely due to the combination of label rarity and our simulation design.
-With few positive labels to begin with, the impact of varying annotator characteristics through our $\beta$ parameter may be overwhelmed by the baseline scarcity of hate speech annotations.
-Ablation Study 2 (Appendix §D) addresses this point.
-Figure 5: Metrics for HS models, by dataset and bias ($\beta$)
-Appendix D Ablation Study 2: Difficult Tweets
-Our simulations assumed that all tweets are impacted the same way (Eq.
-(2)), which is an oversimplification.
-More realistically, annotator characteristics likely have more impact for ambiguous tweets.
-For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$.
-This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq.
-2).
-The filtered datasets contain 267 (OL) and 360 (HS) tweets.
-Figures 6 and 7 show the results for two metrics (ACB, F1) for filtered OL and HS labels.
-In Figure 6(1): the Representative and Adjusted models have similar ACB and are lower than the Non-representative models.
-The F1 scores do not show differences between the models.
-These results are similar to those on the full set of tweets (Figure 3).
-In the two HS figures (7), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics.
-These results are more promising than those on the full set of tweets (Figure 5).
-Figure 6: Metrics for OL models for filtered tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
-Figure 7: Metrics for HS models for filtered tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)